# MQuant: Unleashing the Inference Potential of Multimodal Large Language Models via Full Static Quantization

### v1

@@ 3.2 FHT-Based Weight Outlier Mitigation: Rotation Magnitude Suppression
However, Multi-modal LLMs (MLLMs) combine an LLM component with a visual encoder, which often relies on LayerNorm. Quarot’s partial online transform cannot be applied directly to all such norms. Inspired by SliceGPT (Ashkboos et al., 2024a), we convert visual encoders’ LayerNorms into RMSNorms. More details can be found in the Appendix A.14. This change makes Quarot-like Hadamard transforms applicable to MLLMs. Yet, as Table 2 demonstrates, Quarot still underperforms on many MLLM tasks. Tseng et al. (Tseng et al., 2024) also proved that random Hadamard transforms (RHT) can reduce incoherence for Hessians and weights, but they did not analyze online (fast) Hadamard transforms (FHT). In Quarot, FHT is crucial for low-bit quantization (e.g., 4-bit) as shown in Table 12. We investigate why FHT can yield fresh outliers in MLLMs and propose a method to suppress them.
@@
Figure 3: (a) The pipeline of Quarot, showing offline and partially online Hadamard transforms. (b) An MLLM weight matrix where applying an online FHT produces outliers in the first row.
Channel-Mean Outliers in Online FHT.
@@
Following (Eq. 6), let $W_{\ell_{2}}\in\mathbb{R}^{n\times m}$ be a weight matrix. We denote $\|W_{\ell_{2}}\|_{F}$ as the Frobenius norm. A Hadamard matrix $H\in\{-1,1\}^{n\times n}$ with orthonormal rows preserves this norm, i.e., $\|\,H\,W_{\ell_{2}}\|_{F}=\|W_{\ell_{2}}\|_{F}$. Suppose $\mu_{\ell_{2}}$ is the incoherence coefficient for $W_{\ell_{2}}$, satisfying $\max(|W_{\ell_{2}}|)=\mu_{\ell_{2}}\,\frac{\|W_{\ell_{2}}\|_{F}}{\sqrt{mn}}$. Let $HW_{\ell_{2}}$ be the transformed weight, and let its incoherence coefficient be $\mu_{H\ell_{2}}$. Then
$\frac{\mu_{H\ell_{2}}}{\mu_{\ell_{2}}}\;=\;\frac{\max\bigl{|}\,H\,W_{\ell_{2}}\bigr{|}}{\max\bigl{|}\,W_{\ell_{2}}\bigr{|}}\,.$ (7)
@@
$(HW_{\ell_{2}})_{0j}\;=\;\sqrt{n}\,\mathrm{mean}\bigl{(}w_{:,j}\bigr{)}.$
@@
$\sqrt{n}\,\mathrm{mean}\bigl{(}w_{:,j}\bigr{)}\;>\;\max_{i}\,\bigl{(}w_{ij}\bigr{)},$ (8)
a new channel-mean outlier appears in the first row. Figure 3(b) shows such an occurrence in an MLLM. This issue arises especially in Quarot’s online (partial) FHT step, which applies Hadamard rotations per forward pass rather than strictly offline.
Our Approach: Rotation Magnitude Suppression (RMS).
We propose Rotation Magnitude Suppression (RMS) to handle these new outliers with minimal overhead. We first identify whether a channel meets (Eq. 8). If it does, we:
  1. 1.
    Split the problematic channel from the main GEMM kernel and process it using a separate GEMV.
  2. 2.
    Zero out that row in $HW_{\ell_{2}}$ so that the main kernel does not double-count it.
  3. 3.
    Add the partial output (from the split GEMV) back to the main path before the activation.
Figure 4 shows the workflow. This targeted split ensures large-mean channels do not trigger extreme first-row values during the per-forward-pass FHT. The added cost is small, since only channels meeting Eq. 8 require this procedure.
Figure 4: An overview of our RMS. We separate outlier-prone channels into a dedicated GEMV path, zero their row in the main kernel, and then merge the results.
@@
Summary.
Algorithm 2 outlines how RMS integrates with Quarot’s FHT or other partially online Hadamard transforms. In Table 12, we show that RMS substantially curbs weight outliers in MLLMs under W4A8 quantization, raising accuracy with minimal overhead. Hence, RMS addresses a key shortcoming of online Hadamard transforms by suppressing large-mean channels.
4 Experiments
Table 2: Comprehensive quantization results of different MLLMs across various evaluation datasets.
@@ 4 Experiments
- BF16 BF16 82.82 81.16 782 2153
RTN W8A8 W4A8 7.05 3.70 0.00 140
SQ 9.05 4.10 0.00 148
Quarot 82.00 80.17 782 2115
GLM-4V MQuant 82.06 80.53 782 2164
-9B RTN W4A8 W4A8 7.61 3.60 0.00 163
SQ 9.85 4.40 0.00 188
Quarot 64.16 45.52 516 2048
MQuant 81.58 79.67 754 2120
Models and Datasets.
@@
We evaluate our MQUANT on five MLLMs: InternVL2-8B (Chen et al., 2024a), Qwen-VL-Chat-9.6B (Bai et al., 2023b), MiniCPM-V 2.6-8B (Yao et al., 2024), Qwen2-VL-7B (Wang et al., 2024), and GLM-4V-9B (Hong et al., 2024). Evaluations are conducted on four benchmarks covering OCR and general question answering: TextVQA (Singh et al., 2019), DocVQA (Mathew et al., 2021), OCRBench (Liu et al., 2023), and MME (Fu et al., 2023), which assesses perception and cognition across 14 subtasks. These MLLMs’ details are in Appendix A.11.
@@
Table 3: Comparison of latency and memory saving with Pytorch (BF16), AWQ (W4-only) and ours MQuant (W4A8) on Qwen2-VL-7B. Pytorch and AWQ using the Qwen2-VL-7B official implementation. $\downarrow$ means lower values are better, $\uparrow$ means larger values are better.
@@
Table 4: Multi-Batch speedup comparison of MSQ + AIFS on W4A8 setting. Each row shows the cumulative total of text tokens, images, and textual responses for multi-turn inference.
[…]
Text Image Text bfp16 MQuant bfp16 MQuant bfp16 MQuant
1 10 2240$\times$2240 50 2.54 1.93 +31.6% 18.01 12.89 +39.7% 20.55 14.82 +38.7%
2 10/10 2240$\times$2240 / 2240$\times$2240 50/100 5.42 4.15 +30.6% 37.82 31.56 +19.8% 43.24 35.71 +21.1%
3 10/10/10 2240$\times$2240 / 2240$\times$2240 / 2240$\times$2240 50/100/150 8.24 6.42 +28.3% 48.03 40.35 +19.0% 56.27 46.77 +20.3%
4 10/10/10/10 2240$\times$2240 / 2240$\times$2240 / 2240$\times$2240 / 2240$\times$2240 50/100/150/200 11.17 8.67 +28.9% 59.09 49.92 +8.4% 70.26 58.59 +20.0%
Baselines and Implementation Details.
We test W8A8 and W4A8 quantization settings for both visual encoders and LLMs, comparing RTN, SmoothQuant (Xiao et al., 2022), and Quarot (Ashkboos et al., 2024b). Notably, we apply static per-tensor activation quantization for both components, unlike the dynamic per-token quantization typically used in existing MLLMs. The calibration dataset consists of 256 randomly selected samples from the corresponding benchmark training sets (Singh et al., 2019; Mathew et al., 2021; Liu et al., 2023). The batch size for latency is 1.
4.1 Overall Results
Weight-activation quantization results of various MLLMs.
MQuant can be applied to the quantization of various MLLMs. As shown in Table 2, our MQuant demonstrates significant improvements over several representative quantization methods. In W8A8 setting, MQuant achieves performance nearly equivalent to that of FP models across all evaluation datasets. Notably, even in the more challenging W4A8 setting, MQuant maintains comparable performance with FP models, while other advanced quantization methods exhibit significant performance degradation. These results indicate that our MQuant provide a general and effective PTQ solution with strong compatibility for maintaining high accuracy in MLLMs under various bits settings.
Table 5: Speedup of MSQ + AIFS on W4A8 setting.

### v2

@@ 3.2. FHT-induced Weight Outlier Mitigation: Rotation Magnitude Suppression
Generally, the smaller $\mu$ is, the easier it is to quantize $W$. They also showed that applying a Hadamard transform to both weights and activations can effectively reduce $\mu$. However, Multi-modal LLMs (MLLMs) combine an LLM component with a visual encoder, which often relies on LayerNorm. Quarot’s partial online transform cannot be applied directly to all such norms. Inspired by SliceGPT (Ashkboos et al., 2024a), we convert visual encoders’ LayerNorms into RMSNorms. More details can be found in the Appendix A.14. This change makes Quarot-like Hadamard transforms applicable to MLLMs. Yet, as Table 3 demonstrates, Quarot still underperforms on many MLLM tasks. Quip# (Tseng et al., 2024) also proved that random Hadamard transforms (RHT) can reduce incoherence for Hessians and weights, but they did not analyze online (fast) Hadamard transforms (FHT). In Quarot, FHT is crucial for low-bit quantization (e.g., 4-bit) as shown in Appendix  A.9. We investigate why FHT can yield fresh outliers in MLLMs and propose a method to suppress them. Besides, due to the difference of MLLMs architecture, we also discuss the rotation-based implementation details before RMS in Appendix  A.13.
@@
Figure 4. (a) The pipeline of Quarot, showing offline and partially online Fast Hadamard transforms. (b) Weight matrix where applying online FHT produces outliers in the first row.
Channel-Mean Outliers in Online FHT
@@
Following (Eq. 6), let $W_{\ell_{2}}\in\mathbb{R}^{n\times m}$ be a weight matrix. A Hadamard matrix $H\in\{-1,1\}^{n\times n}$ with orthonormal rows preserves this norm, i.e., $\|\,H\,W_{\ell_{2}}\|_{F}\\ =\|W_{\ell_{2}}\|_{F}$. Suppose $\mu_{\ell_{2}}$ is the incoherence coefficient for $W_{\ell_{2}}$, satisfying $\max(|W_{\ell_{2}}|)=\mu_{\ell_{2}}\,\frac{\|W_{\ell_{2}}\|_{F}}{\sqrt{mn}}$. Let $HW_{\ell_{2}}$ be the transformed weight, and let its incoherence coefficient be $\mu_{H\ell_{2}}$. Then we have:
(7) $\frac{\mu_{H\ell_{2}}}{\mu_{\ell_{2}}}\;=\;\frac{\max\bigl{|}\,H\,W_{\ell_{2}}\bigr{|}}{\max\bigl{|}\,W_{\ell_{2}}\bigr{|}}\,.$
@@
(8) $(HW_{\ell_{2}})_{0j}\;=\;\sqrt{n}\,\mathrm{mean}\bigl{(}w_{:,j}\bigr{)}.$
@@
(9) $\sqrt{n}\,\mathrm{mean}\bigl{(}w_{:,j}\bigr{)}\;>\;\max_{i}\,\bigl{(}w_{ij}\bigr{)},$
a new channel-mean outlier appears in the first row. Figure 4(b) shows such an occurrence in an MLLM. This issue arises especially in Quarot’s online (partial) FHT step, which applies Hadamard rotations per forward pass rather than strictly offline. The detailed approve in Appendix  A.3.
Table 2. Compliance Ratio with Eq. 9 Across Model Components
Model Part All Blocks Meet Eq. 9 Ratio
Number Number
Internvl2-8B Visual 24 24 100%
LLM 32 4 12.5%
Qwenvl-9.6B Visual 48 48 100%
LLM 32 13 40.6%
MiniCPM-V-2.6-8B Visual 27 27 100%
LLM 28 3 10.7%
GLM-4V-9B Visual 63 63 100%
LLM 40 6 15.0%
Qwen2vl-7B Visual 32 32 100%
LLM 28 1 3.5%
Prevalence of Eq. 9 Compliance
We conduct systematic block-level verification across five state-of-the-art MLLMs (InternVL2-8B, QwenVL-9.6B, MiniCPM-V-2.6-8B, GLM-4V-9B, and Qwen2VL-7B) to quantify the occurrence of Fast Hadamard Transform (FHT)-induced weight outliers. As evidenced in table  2, our analysis reveals two critical findings: (1) Universal Presence in Vision Encoders: All visual $down_{proj}$ layers (100%) exhibit FHT-induced weight outliers that satisfy Eq. 9, as conceptually illustrated in Fig. 4(b). (2) Variable Manifestation in LLMs: The phenomenon appears in only 3%–40% of LLM $down_{proj}$ layers across different architectures. This empirical validation confirms above theoretical soundness and the critical need for specialized handling of FHT-generated weight outliers, particularly in vision components where their universal presence fundamentally impacts model quantization.
Rotation Mitigation Scheme (RMS)
To this end, we propose RMS to handle these new outliers with minimal overhead. We first identify whether a channel meets (Eq. 9). If it does, we: Split the problematic channel from the main GEMM kernel and process it using a separate GEMV. Zero out that row in $HW_{\ell_{2}}$ so that the main kernel does not double-count it. Add the partial output (from the split GEMV) back to the main path before the activation. Figure 5 shows the workflow. This targeted split ensures large-mean channels do not trigger extreme first-row values during the per-forward-pass FHT. The added cost is small, since only channels meeting Eq. 9 require this procedure. Detailed RMS algorithm in Appendix A.7 outlines how RMS integrates with Quarot’s FHT. Overall, RMS addresses a key shortcoming of online Hadamard transforms by suppressing large-mean channels.
Figure 5. An overview of our proposed RMS. We separate outlier-prone channels into a dedicated GEMV path, zero their row in the main kernel, and then merge the results.
@@
4. Experiments
Table 3. Comprehensive quantization results of different MLLMs across various multimodal reasoning datasets.
@@ 4. Experiments
- BF16 BF16 85.48 95.95 883 2479
RTN W8A8 W4A8 37.21 40.94 426 1444
SQ 50.33 55.41 480 1601
Quarot 80.03 91.21 781 2299
Qwen2-VL MQuant 85.48 95.90 880 2469
-72B RTN W4A8 W4A8 28.21 25.94 426 1137
SQ 47.11 54.95 413 1586
Quarot 71.86 86.11 701 2264
MQuant 85.03 95.49 868 2471
Models and Datasets
@@
We evaluate our MQUANT on five MLLMs: InternVL2-8B (Chen et al., 2024b), Qwen-VL-Chat-9.6B (Bai et al., 2023b), MiniCPM-V 2.6-8B (Yao et al., 2024), Qwen2-VL-7B (Wang et al., 2024a), and GLM-4V-9B (Hong et al., 2024). Evaluations are conducted on four benchmarks covering OCR and general question answering: TextVQA (Singh et al., 2019), DocVQA (Mathew et al., 2021), OCRBench (Liu et al., 2023), and MME (Fu et al., 2023), which assesses perception and cognition across 14 tasks. These MLLMs’ details are in Appendix A.11.
Baselines and Implementation Details
We test W8A8 and W4A8 quantization settings for both visual encoders and LLMs, comparing RTN, SmoothQuant (Xiao et al., 2022), and Quarot (Ashkboos et al., 2024b). Notably, we apply static per-tensor activation quantization for both components, unlike the dynamic per-token quantization typically used in existing MLLMs. The calibration dataset consists of 256 randomly selected samples from the corresponding benchmark training sets (Singh et al., 2019; Mathew et al., 2021; Liu et al., 2023). The batch size for latency measurement is 1.
4.1. Overall Results
As shown in Table 3, MQuant can apply to the quantization of various MLLMs, and demonstrates significant improvements over several representative quantization methods. In W8A8 setting, MQuant achieves near-lossless performance to the FP models across diverse datasets and different model scales (7-72B). Notably, even in the more challenging W4A8 setting, MQuant maintains comparable performance with FP models, while other advanced quantization methods exhibit significant performance degradation. These results indicate that our MQuant provide a general and effective PTQ solution with strong compatibility for maintaining high accuracy in MLLMs under various bits settings across diverse multi-modal reasoning tasks. This suggests MQuant maintains robustness (e.g., stability to typical input or models changes) in these frameworks.
Table 4. Comparison of existing MLLMs Quantization Methods
Method W bits A bits Scale GEMM kernel Extra ops
Q-VLM (Wang et al., 2024b) NF4 INT4 per-channel static BF16 GEMM dequant W&A
QSLAW (Xie et al., 2024) INT4 BF16 weight-only BF16 GEMM dequant W
MBQ (Li et al., 2024a) INT4 INT4 per-token dynamic INT4 GEMM per token scale
MQuant (ours) INT4 INT4 per-modality static INT4 GEMM none
Table 5. ScienceQA (Lu et al., 2022) test Results on LLaVA1.5-13B
[…]
BF16 - 90.19 93.14 87.09 89.39 87.06 89.83 90.00
w4a4 Q-VLM (Wang et al., 2024b) 82.55 73.32 83.18 81.03 70.82 86.74 80.78
w4a4 MQuant 87.07 85.24 87.05 88.49 85.50 89.77 87.18
w4 only QSLAW (Xie et al., 2024) 83.26 91.79 80.00 82.80 81.30 83.07 83.70
w4 only MQuant (g=128) 90.21 91.95 87.05 89.20 86.66 89.94 89.16
Table 6. Performance Comparison under Different Precision and Methods (T.VQA, D.VQA, OCRB, MME) on Qwen2VL-72B.
[…]
BF16 - 85.48 95.95 883 2479
W4A8 MBQ (Li et al., 2024a) 83.41 92.30 854 2402
W4A8 MQuant (ours) 85.03 95.49 868 2471
4.2. Comparison with other MLLMs Quantization Methods
To rigorously evaluate MQuant, we compare it against recent state-of-the-art (SOTA) quantization methods (Wang et al., 2024b; Xie et al., 2024; Li et al., 2024a) specifically designed for MLLMs. Table 4 provides a high-level comparison of the quantization schemes. Existing methods often introduce substantial runtime overhead. For example, MBQ (Li et al., 2024a) employs online dynamic per-token quantization; Q-VLM (Wang et al., 2024b) and QSLAW (Xie et al., 2024) require dequantization and perform GEMM in resource-intensive BF16 precision. In contrast, MQuant is the only method that achieves a fully static, per-modality quantization scheme that is directly compatible with highly efficient INT4 GEMM kernels. This design choice is the foundation of its superior inference speed. Besides, due to their different evaluation benchmark, we compare them on their reported results. The empirical results in Table 5 and Table 6 confirm MQuant’s accuracy advantage. On the ScienceQA benchmark (Lu et al., 2022) (Table 5) in LLaVA1.5-13B (Liu et al., 2024), our W4A4 MQuant surpasses Q-VLM by a significant margin of 6.4% in average accuracy. In the weight-only setting, MQuant outperforms QSLAW by 5.46%, highlighting the effectiveness of our activation quantization strategy. Similarly, on Qwen2VL-72B(Wang et al., 2024a) (Table 6), our W4A8 MQuant achieves near-lossless performance, outperforming MBQ while avoiding its costly per-token dynamic quantization. This superior performance is a direct result of our novel designs. The MSQ effectively handles the distributional disparity between visual and textual tokens, a challenge that other methods do not explicitly address. Furthermore, the AIFS scheme eliminates the need for dynamic scales without compromising model accuracy, enabling the use of fast, hardware-friendly static quantization. Together with RMS, these contributions allow MQuant to achieve both the highest accuracy and the fastest true INT4 inference among existing MLLM quantization methods.
4.3. Ablation Study
In this section, we select Qwen2-VL-7B (Wang et al., 2024a), one of the most powerful open-source MLLMs, to validate the effectiveness of our designs.
Table 7. Ablation study of proposed designs on Qwen2-VL-7B (Wang et al., 2024a) with W4A8 setting.
[…]
Static AIFS + MSQ RMS
BF16 84.43 93.87 842 2319 6523
71.44 83.96 670 1911 5479
78.95 87.55 721 2095 5484
84.32 93.58 824 2255 5471
Ablation Study of Proposed Quantization Designs
We perform ablations on Qwen2-VL-7B (Table 7) to isolate the contributions of each quantization component. Beginning with a baseline that applies GPTQ + Hadamard transformations to both the LLM and vision parts with per-tensor static quantization, we progressively introduce MSQ+AIFS for the LLM. This significantly boosts accuracy and latency, highlighting their effectiveness on multimodal input tokens. Next, incorporating RMS to suppress newly arising outliers achieves near-floating-point performance, underscoring the overall robustness and efficiency of our pipeline. We also demonstrate that the design of RMSis also effective for LLMs in Appendix  A.5.
Table 8. Ablation study about calibration dataset size.
[…]
BF16 84.43 93.87 842 2319
128 84.28 93.50 820 2243
256 84.32 93.58 824 2255
512 84.32 93.57 823 2254
Selection of Calibration Dataset Size
Following  (Xiao et al., 2022; Ashkboos et al., 2024b), we selected a sample subset to get the statistics of activations during the calibration process. This ensures the calibration data reflects the distribution in each dataset. To assess sensitivity to calibration dataset size of MQuant, we conducted experiments on Qwen2-VL-7B, including [128, 256, 512] samples. The Table 8 shows the consistency of MQuant in W4A8 with loessless accuracy. These results, revised in supplementary, demonstrate that MQuant is largely insensitive to calibration dataset size. We believe this robustness enhances MQuant’s practical applicability. Therefore, we select 256 randomly selected samples from the corresponding benchmark training set.
Table 9. Latency combining AIFS with FlashAttention from 100 to 8100 tokens ($Q/K/V$ shape is (1, 28, num-tokens, 128))
Tokens Number 100 900 2500 4900 8100
FlashAttn (ms) 27.3 33.6 35.5 165.9 271.6
AIFS+FlashAttn (ms) 27.3 34.1 36.4 167.5 272.1
Table 10. The accuracy and speedup of our MSQ and AIFS scheme on the linear layer during prefill stage. The input sequence is structured as "text-image-text-image-text-image-text" with an image resolution of $2240\times 2240$ and 50 textual tokens. Latency were tested on an NVIDIA RTX 6000 Ada Generation.
[…]
BF16 BF16 84.43 93.87 842 2319 1.690 -
BF16 W4-g128(AWQ) 83.93 (-0.50) 93.13 (-0.74) 828 (-14) 2252 (-67) 2.057 (+0.367) -17.8%
@@ 4.3. Ablation Study
Table 11. Comparison of latency and memory saving with Pytorch (BF16), AWQ (W4-only) and MQuant (W4A8) on Qwen2-VL-7B. $\downarrow$ means lower values are better, $\uparrow$ means larger values are better.
@@
Integrate AIFS with Flash Attention
As evidenced in Table 9, we integrated the causal attention mechanism of AIFS into FlashAttention and tested it on an RTX 6000, observing that its speed is nearly identical to that of standard causal attention. The results demonstrates seamless integration MBQ+AIFS with Flash Attention, exhibiting negligible latency overhead (<0.3% across all tested token lengths). The minimal performance gap (maximum 1.6 ms) stems from two key design choices: (1) AIFS’s token reordering operates exclusively in the preprocessing phase without interfering with core attention computation, and (2) our metadata propagation mechanism (visual token indices m/n) preserves Flash Attention’s native memory access patterns. Notably, the 0.96% latency increase at 4,900 tokens validates strong scalability for high-resolution inputs. This highlight the hardware compliance advantages of MBQ and AIFS designs: CUDA-aligned token layouts enable binary compatibility with FlashAttention kernels, which is particularly effective for real-time processing of high-resolution multimodal inputs.
Figure 6. Illustration the weight distributions for the down-proj layer (block 21) in Qwen2VL-7B’s visual encoder under three conditions: (a) original weight, (b) weights after FHT, and (c) weights after FHT with our RMS. More resutls refer to Appendix A.15
Weight Distribution with Rotation Magnitude Suppression (RMS)
As shown in Fig.  12, we visualizes the magnitude of weights outliers without (w/o) and with (w/) RMS. The weight distribution reveals that FHT amplification increases weight magnitudes, while RMS effectively reduces these magnitudes by a factor of 1.0 to 0.1, significantly improving quantization stability (Other distribution are in Appendix A.15). This is also consistent with our ablation results, shown that RMS significantly enhances performance for both MLLMs (Table  7). Taking Qwen2-VL as example, the vision encoder’s down-projection layer (3584×1280, C_inC_outC\_in*C\_out) and LLM’s corresponding layer (18944×3584) demonstrate the efficiency of our approach. The RMS implementation converts a W4A8 GEMM operation into a combined W4A8 GEMM + W4A8 GEMV computation, where the GEMV is selectively applied only to affected channels. This introduces negligible overhead, precisely 1/3584 and 1/18944 of the total computation for vision and LLMs parts.
Figure 7. The Speedup of AIFS+MSQ on Qwen2-VL-7B.
Accuracy & Speedup of AIFS
4.4. Latency and Memory Analysis
We evaluate MQuant’s performance across three key metrics: (1) inference speedup, (2) memory efficiency (tested over image resolutions from $280^{2}$-$5600^{2}$), and (3) decoding acceleration. All experiments employ the established ”text-image-text” input format in (Duan et al., 2024) with a fixed 15 textual tokens unless explicitly varied.
Table 12. Speedup of MSQ + AIFS on W4A8 setting.